# When Ads Become Profiles: Uncovering the Invisible Risk of Web Advertising at Scale with LLMs

### v2

@@ Abstract.
Regulatory limits on explicit targeting have not eliminated algorithmic profiling on the Web, as optimisation systems still adapt ad delivery to users’ private attributes. The widespread availability of powerful zero-shot multimodal Large Language Models (LLMs) has dramatically lowered the barrier for exploiting these latent signals for adversarial inference. We investigate this emerging societal risk, specifically how adversaries can now exploit these signals to reverse-engineer private attributes from ad exposure alone. We introduce a novel pipeline that leverages LLMs as adversarial inference engines to perform natural language profiling. Applying this method to a longitudinal dataset comprising over 435,000 ad impressions collected from 891 users, we conducted a large-scale study to assess the feasibility and precision of inferring private attributes from passive online ad observations. Our results demonstrate that off-the-shelf LLMs can accurately reconstruct complex user private attributes, including party preference, employment status, and education level, consistently outperforming strong census-based priors and matching or exceeding human social perception, while operating at only a fraction of the cost (223× lower) and time (52× faster) required by humans. Critically, actionable profiling is feasible even within short observation windows, indicating that prolonged tracking is not a prerequisite for a successful attack. These findings provide the first empirical evidence that ad streams serve as a high-fidelity digital footprint, enabling off-platform profiling that inherently bypasses current platform safeguards, highlighting a systemic vulnerability in the ad ecosystem and the urgent need for responsible web AI governance in the generative AI era. The code is available at https://github.com/Breezelled/when-ads-become-profiles.
Targeted Advertising, Large Language Models, Privacy Risks
copyright: noneconference: ; ;
@@ 1. Introduction
Online advertising serves as the primary economic engine of the Web, enabling the free distribution of content and services. To maximise relevance and revenue, platforms like Facebook and Google deploy sophisticated, opaque algorithms to deliver personalised content, matching advertisements to users based on granular inferred profiles and behaviours (Computational_Advertising_Recent_Advances). However, as privacy concerns have mounted, the Web ecosystem has shifted toward stricter governance and regulatory compliance (GDPR2016). A pivotal moment occurred in 2022, when major platforms like Meta removed ”Detailed Targeting” options for sensitive categories, including political beliefs, health causes, sexual orientation, and religious practices, to prevent potential abuse and discrimination  [^1] https://www.facebook.com/business/news/removing-certain-ad-targeting-options-and-expanding-our-ad-controls. This shift was intended to enhance user privacy by limiting how advertisers could directly target vulnerable groups through the platform’s native tools.
@@
While these policies limit explicit targeting options, they do not neutralise the demographic skews introduced by the platform’s optimisation algorithms. Prior studies have established that ad delivery remains highly correlated with private attributes, even under neutral targeting settings (ali2019discrimination; imana2021job; ali2022all; ali2023problematic), as an artifact of algorithmic optimisation for user engagement. Our own preliminary analysis confirms that these correlations persist within the dataset we use, creating a discernible link between ad exposure and user private attributes. Consequently, the stream of advertisements shown to a user continues to serve as a high-fidelity proxy for their identity, containing latent fine-grained signals about their private attributes.
@@
The emergence of generative AI challenges the efficacy of existing governance frameworks, introducing potential vulnerabilities by diversifying attack pathways and increasing uncertainty. Previously, exploiting these signals required significant resources: collecting massive labeled datasets and training classifiers (bi2013inferring; kosinski2013private; schwartz2013personality; hinds2018demographic; tricomi2024spotify). However, this barrier has been fundamentally altered by the rapid democratisation of powerful generative AI like Large Language Models (LLMs). Such technologies demonstrate remarkable abilities to understand and generate nuanced, human-like content (hurst2024gpt; openai2025gpt5systemcard), reason about complex contexts (jaech2024openai; guo2025deepseek), adapt to user-specific preferences (zhang2025personalization; openai2025memory), and even integrate multimodal information (hurst2024gpt; openai2025gpt5systemcard; anthropic2025claude4; comanici2025gemini), bringing immense convenience but also raising new societal concerns in domains such as online advertising (AI_Driven_Online_Advertising_Market_Design_Generative_AI_and_Ethics). As these models become increasingly accessible and capable (Computational_Advertising_Recent_Advances), not only through APIs or open-source implementations but also via free, public-facing web interfaces to some of the most advanced proprietary models (comanici2025gemini; hurst2024gpt; openai2025gpt5systemcard; anthropic2025claude4; guo2025deepseek), which dramatically lower the barrier for conducting sophisticated analyses at scale, including those with potentially harmful intent.
@@
Our findings reveal that LLMs can reconstruct private attributes such as age, gender, education, employment status and party preference with high accuracy, outperforming strong census-based priors and matching human annotators. Crucially, we demonstrate that this inference is feasible even with short observation windows (session-level), meaning prolonged tracking is not a prerequisite for a successful attack. These findings provide the first empirical demonstration that multimodal LLMs can reverse-engineer user private attributes from passive observation alone, highlighting the scale of privacy risks and the urgent need for safeguards in the era of accessible, high-capacity generative AI.
@@ 2.1. Digital Footprints and Traditional Inference
Prior to the emergence of LLMs, research had established that digital footprints are predictive of private user attributes. Studies have demonstrated that diverse traces, ranging from Facebook Likes (kosinski2013private) and social media language use (schwartz2013personality) to search logs (bi2013inferring) and even public Spotify playlists (tricomi2024spotify), contain latent demographic signals. In the visual domain, tasks such as human attribute recognition and pedestrian attribute recognition have achieved high precision in identifying traits like gender and age from images (app10165608; WANG2022108220). However, these traditional approaches typically required significant resources, relying on massive labeled datasets to train bespoke classifiers via supervised learning (hinds2018demographic).
Within the advertising domain, extensive auditing literature has similarly documented systematic demographic skews in ad delivery algorithms (ali2019discrimination; imana2021job; ali2022all; ali2023problematic). While these studies originally aimed to detect algorithmic discrimination, they provide the crucial empirical evidence that ad streams carry latent signals. Our work builds upon this foundation by shifting the lens to adversarial inference, demonstrating how these established correlations can be exploited by off-the-shelf LLMs without the need for the resource-intensive training pipelines required by previous methods.
@@ 2.2. Utility and Risks of LLM Profiling
The rise of LLMs has fundamentally altered user profiling, shifting from opaque vector representations to transparent, language-based reasoning. In recommender systems, this capability is widely leveraged to construct scrutable user models from interaction histories (balog2019transparent; radlinski2022natural; ramos-etal-2024-transparent; zhou2024language), or to infer detailed preferences from sparse inputs like POI check-ins and collaborative signals (wongso2024genup; ren2024representation; xi2024towards). Beyond recommendation, LLMs demonstrate strong reasoning capabilities in broader user modeling tasks, ranging from inferring latent interests to generating personalized content (tan2023user; meguellati2024how).
Recently, this inferential power has raised significant concerns regarding inferential privacy, where adversaries exploit models to deduce private information from user content. Studies have demonstrated that LLMs can infer personal attributes from online text posts (staab24beyond), highlighting that such automated inference operates at a fraction of the cost and time required by human analysts. Extending this to the visual domain, recent works utilise vision-language models to extract attributes from user-uploaded images (tomekcce2024private) and aggregate clues across personal photo albums using agentic frameworks (liu2025eye). Our work identifies a critical gap in this emerging literature. While previous studies focus on active digital footprints, content users voluntarily create or upload, we investigate passive digital footprints: the stream of advertisements pushed to users by algorithmic systems. By demonstrating that off-the-shelf LLMs can reverse-engineer user private attributes from these passive exposures, we uncover a systemic vulnerability where the platform’s own targeting logic can be exploited to compromise user privacy.
@@ 3.1. Adversary Goal and Capabilities
The adversary aims to infer a target user’s private attributes $\mathcal{A}$ (e.g., party preference, employment status) by observing the stream of advertisements $\mathcal{S}$ delivered to them. This adversary seeks to exploit the real-time, high-fidelity targeting intelligence embedded in the ad platform itself. This approach differs fundamentally from existing data acquisition methods. Unlike scraping public social media profiles (staab24beyond; tomekcce2024private), which relies on active user disclosure, this attack exploits passive exposure. Furthermore, unlike purchasing data from brokers and targeting from the platform, which creates discoverable records and is subject to regulatory scrutiny.
@@ 3.2. Attack Vector
We identify browser extensions that abuse legitimate privileges as the potential primary vector for this attack. This scenario is particularly severe due to its inherent stealth and scalability. Rather than distributing specialised malware, an adversary can opportunistically deploy this attack within the existing ecosystem of widely installed, benign-functioning extensions, such as ad blockers, coupon finders, or page translators. These extensions legitimately require permissions to read web page content to function, providing a perfect cover for data harvesting (singh2025study).
This vector effectively bypasses both user and platform scrutiny. Users typically focus their privacy concerns on invisible trackers and cookies while overlooking the semantic signals embedded in visible ad creatives (yao2017folk). Similarly, app or plugin store reviews often focus on code safety rather than the privacy implications of data inferred from legitimate content access (singh2025study). This creates a regulatory blind spot where a benign-looking extension can silently harvest ad content without triggering security alarms. Furthermore, by leveraging the zero-shot capabilities of LLMs, the adversary can automate this process for scalable profiling. As we demonstrate in Section 5.3.1, actionable profiles can be reconstructed from short observation windows, meaning the adversary does not need to maintain long-term persistence to be effective. Ultimately, this vector offers a distinct strategic advantage: it enables off-platform profiling that evades the platform’s own privacy safeguards, such as the removal of sensitive targeting options. The adversary effectively exploit the optimisation logic of the ad platform to build sensitive user profiles at a low marginal cost and without leaving an audit trail.
@@ 4.1. Multimodal Ad Understanding
To transform raw, multimodal ad data into a structured representation suitable for profiling, we employ an LLM-driven extraction process. Let each advertisement $d_{i}$ in our dataset consist of visual content $V_{i}$ (one or more images) and textual metadata $T_{i}$ (title, body, call-to-action). We utilise Gemini 2.0 Flash (gemini2_flash_modelcard) as our inference engine $\mathcal{M}$, selected for its balance of multimodal reasoning and cost-efficiency required to process $N>435,000$ impressions. For each ad, conditioned on an extraction prompt $\mathcal{P}_{\text{extract}}$, the model $\mathcal{M}$ maps the input into a structured feature set $F_{i}$:
(1) $F_{i}=\mathcal{M}(V_{i},T_{i}\mid\mathcal{P}_{\text{extract}})=\{c_{i},\mathcal{D}_{i},\mathcal{L}_{i},\mathcal{E}_{i}\}$
This structured output $F_{i}$ encapsulates four semantic components: (1) a Caption ($c_{i}$) summarising the core message and visual elements of the ad; (2) Descriptive Categories ($\mathcal{D}_{i}$), a set of free-form labels characterising tone, style, and communication strategy; (3) IAB Categories ($\mathcal{L}_{i}$), a multi-label zero-shot classification into a controlled vocabulary of 45 options based on the IAB taxonomy (iab_ad_product_taxonomy); and (4) Key Entities ($\mathcal{E}_{i}$), a list of extracted specific brands, products, or locations. This transformation $d_{i}\to F_{i}$ extracts high-dimensional multimodal signals into a concise textual representation. We validate the quality of these model-generated features in Appendix C and discuss a case of Human–AI disagreement in Appendix D.
@@ 4.2. Session-Level User Profile Reconstruction
To capture short-term profiling signals, we segment the continuous user stream $\mathcal{S}_{u}$ into discrete sessions. Let a session $\mathcal{S}_{s}$ be a chronologically ordered sequence of $n$ structured ad features observed within a continuous browsing window (segmentation details in Section 5.2):
(2) $\mathcal{S}_{s}=(F_{s_{1}},F_{s_{2}},\dots,F_{s_{n}})$
where each $F_{s_{i}}$ corresponds to the structured feature set extracted in the previous stage. We construct a textual representation $T_{s}$ by concatenating the components of each $F\in\mathcal{S}_{s}$ in temporal order.
This text-centric transformation is a deliberate methodological choice designed to mitigate two critical limitations of processing raw visual sequences. First, computational feasibility: user sessions vary significantly in length ($n\in[3,50]$). Processing such long sequences of high-resolution images exceeds the token limits and budget constraints of current API-based models. Second, contextual fidelity: presenting a collage of disparate ad images simultaneously can induce visual-textual alignment ambiguity, where the model struggles to associate specific textual claims with their corresponding visual creatives. We define a session-level inference function that maps the textual representation $T_{s}$, conditioned on a profiling prompt $\mathcal{P}_{\text{profile session}}$, to a prediction of user’s private attributes $\hat{\mathcal{A}}_{s}$ and a reasoning summary $r_{s}$:
@@
Here, $\hat{\mathcal{A}}_{s}$ contains zero-shot classifications for the six private attributes derived from the current session. The model generates a reasoning summary ($r_{s}$), a natural language explanation contains the key demographic signals identified within that specific session. This summary serves as the compact information carrier for the subsequent longitudinal analysis.

### v3

@@ Abstract.
Regulatory limits on explicit targeting have not eliminated algorithmic profiling on the Web, as optimisation systems still adapt ad delivery to users’ private attributes. The widespread availability of powerful zero-shot multimodal Large Language Models (LLMs) has dramatically lowered the barrier for exploiting these latent signals for adversarial inference. We investigate this emerging societal risk, specifically how adversaries can now exploit these signals to reverse-engineer private attributes from ad exposure alone. We introduce a novel pipeline that leverages LLMs as adversarial inference engines to perform natural language profiling. Applying this method to a longitudinal dataset comprising over 435,000 Facebook ad impressions collected from 891 users, we conducted a large-scale study to assess the feasibility and precision of inferring private attributes from passive online ad observations. Our results demonstrate that off-the-shelf LLMs can accurately reconstruct complex user private attributes, including party preference, employment status, and education level, consistently outperforming strong census-based priors and matching or exceeding human social perception at only a fraction of the cost (223× lower) and time (52× faster) required by humans. Critically, actionable profiling is feasible even within short observation windows, indicating that prolonged tracking is not a prerequisite for a successful attack. These findings provide the first empirical evidence that ad streams serve as a high-fidelity digital footprint, enabling off-platform profiling that inherently bypasses current platform safeguards, highlighting a systemic vulnerability in the ad ecosystem and the urgent need for responsible web AI governance in the generative AI era. The code is available at https://github.com/Breezelled/when-ads-become-profiles.
Targeted Advertising; Large Language Models; Privacy Risks
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3793060isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Online advertisingccs: Security and privacy Social aspects of security and privacyccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Online advertising serves as the primary economic engine of the Web, enabling the free distribution of content and services. To maximise relevance and revenue, platforms like Facebook and Google deploy sophisticated, opaque algorithms to deliver personalised content, matching advertisements to users based on granular inferred profiles and behaviours (Pan et al., 2025). However, as privacy concerns have mounted, the Web ecosystem has shifted toward stricter governance and regulatory compliance (European Parliament and Council of the European Union, 2016). A pivotal moment occurred in 2022, when major platforms like Meta removed ”Detailed Targeting” options for sensitive categories, including political beliefs, health causes, sexual orientation, and religious practices, to prevent potential abuse and discrimination (Meta, 2022). This shift was intended to enhance user privacy by limiting how advertisers could directly target vulnerable groups through the platform’s native tools.
@@
While these policies limit explicit targeting options, they do not neutralise the demographic skews introduced by the platform’s optimisation algorithms. Prior studies have established that ad delivery remains highly correlated with private attributes, even under neutral targeting settings (Ali et al., 2019; Imana et al., 2021; Ali et al., 2022, 2023), as an artifact of algorithmic optimisation for user engagement. Our own preliminary analysis confirms that these correlations persist within the dataset we use, creating a discernible link between ad exposure and user private attributes. Consequently, the stream of advertisements shown to a user continues to serve as a high-fidelity proxy for their identity, containing latent fine-grained signals about their private attributes.
@@
The emergence of generative AI challenges the efficacy of existing governance frameworks, introducing potential vulnerabilities by diversifying attack pathways and increasing uncertainty. Previously, exploiting these signals required significant resources: collecting massive labeled datasets and training classifiers (Bi et al., 2013; Kosinski et al., 2013; Schwartz et al., 2013; Hinds and Joinson, 2018; Tricomi et al., 2024). However, this barrier has been fundamentally altered by the rapid democratisation of powerful generative AI like Large Language Models (LLMs). Such technologies demonstrate remarkable abilities to understand and generate nuanced, human-like content (Hurst et al., 2024; OpenAI, 2025), reason about complex contexts (Jaech et al., 2024; Guo et al., 2025), adapt to user-specific preferences (Zhang et al., 2025; OpenAI, 2024), and even integrate multimodal information (Hurst et al., 2024; OpenAI, 2025; Anthropic, 2025; Comanici et al., 2025), bringing immense convenience but also raising new societal concerns in domains such as online advertising (He et al., 2024). As these models become increasingly accessible and capable (Pan et al., 2025), not only through APIs or open-source implementations but also via free, public-facing web interfaces to some of the most advanced proprietary models (Comanici et al., 2025; Hurst et al., 2024; OpenAI, 2025; Anthropic, 2025; Guo et al., 2025), which dramatically lower the barrier for conducting sophisticated analyses at scale, including those with potentially harmful intent.
@@
Our findings reveal that LLMs can reconstruct private attributes such as age, gender, education, employment status and party preference with high accuracy, outperforming strong census-based priors and matching human annotators. Crucially, we demonstrate that this inference is feasible even with short observation windows (session-level), meaning prolonged tracking is not a prerequisite for a successful attack. These findings provide the first empirical demonstration that multimodal LLMs can reverse-engineer user private attributes from passive observation (passive observation here means passively viewed ads from the user’s perspective) alone, highlighting the scale of privacy risks and the urgent need for safeguards in the era of accessible, high-capacity generative AI.
@@ 2.1. Digital Footprints and Traditional Inference
Prior to the emergence of LLMs, research had established that digital footprints are predictive of private user attributes. Studies have demonstrated that diverse traces, ranging from Facebook Likes (Kosinski et al., 2013) and social media language use (Schwartz et al., 2013) to search logs (Bi et al., 2013) and even public Spotify playlists (Tricomi et al., 2024), contain latent demographic signals. In the visual domain, tasks such as human attribute recognition and pedestrian attribute recognition have achieved high precision in identifying traits like gender and age from images (Yaghoubi et al., 2020; Wang et al., 2022). However, these traditional approaches typically required significant resources, relying on massive labeled datasets to train bespoke classifiers via supervised learning (Hinds and Joinson, 2018).
Within the advertising domain, extensive auditing literature has similarly documented systematic demographic skews in ad delivery algorithms (Ali et al., 2019; Imana et al., 2021; Ali et al., 2022, 2023). While these studies originally aimed to detect algorithmic discrimination, they provide the crucial empirical evidence that ad streams carry latent signals. Our work builds upon this foundation by shifting the lens to adversarial inference, demonstrating how these established correlations can be exploited by off-the-shelf LLMs without the need for the resource-intensive training pipelines required by previous methods.
@@ 2.2. Utility and Risks of LLM Profiling
The rise of LLMs has fundamentally altered user profiling, shifting from opaque vector representations to transparent, language-based reasoning. In recommender systems, this capability is widely leveraged to construct scrutable user models from interaction histories (Balog et al., 2019; Radlinski et al., 2022; Ramos et al., 2024; Zhou et al., 2024), or to infer detailed preferences from sparse inputs like POI check-ins and collaborative signals (Wongso et al., 2025; Ren et al., 2024; Xi et al., 2024). Beyond recommendation, LLMs demonstrate strong reasoning capabilities in broader user modeling tasks, ranging from inferring latent interests to generating personalized content (Tan and Jiang, 2023; Meguellati et al., 2024).
Recently, this inferential power has raised significant concerns regarding inferential privacy, where adversaries exploit models to deduce private information from user content. Studies have demonstrated that LLMs can infer personal attributes from online text posts (Staab et al., 2024), highlighting that such automated inference operates at a fraction of the cost and time required by human analysts. Extending this to the visual domain, recent works utilise vision-language models to extract attributes from user-uploaded images (Tömekçe et al., 2024) and aggregate clues across personal photo albums using agentic frameworks (Liu et al., 2025). Our work identifies a critical gap in this emerging literature. While previous studies focus on active digital footprints, content users voluntarily create or upload, we investigate passive digital footprints: the stream of advertisements pushed to users by algorithmic systems. By demonstrating that off-the-shelf LLMs can reverse-engineer user private attributes from these passive exposures, we uncover a systemic vulnerability where the platform’s own targeting logic can be exploited to compromise user privacy.
@@ 3.1. Adversary Goal and Capabilities
The adversary aims to infer a target user’s private attributes $\mathcal{A}$ (e.g., party preference, employment status) by observing the stream of advertisements $\mathcal{S}$ delivered to them. This adversary seeks to exploit the real-time, high-fidelity targeting intelligence embedded in the ad platform itself. This approach differs fundamentally from existing data acquisition methods. Unlike scraping public social media profiles (Staab et al., 2024; Tömekçe et al., 2024), which relies on active user disclosure, this attack exploits passive exposure. Furthermore, purchasing data from brokers and targeting from the platform create discoverable records and are subject to regulatory scrutiny.
@@ 3.2. Attack Vector
We identify browser extensions that abuse legitimate privileges as the potential primary vector for this attack. This scenario is severe due to its inherent stealth and scalability. Rather than distributing specialised malware, an adversary can opportunistically deploy this attack within the existing ecosystem of widely installed, benign-functioning extensions, such as ad blockers, coupon finders, or page translators. These extensions legitimately require permissions to read web page content to function, providing a perfect cover for data harvesting (Singh et al., 2025). In contrast, collecting URL histories or full-page logs is more likely to be flagged by extension reviews.
This vector effectively bypasses user and platform scrutiny. Users typically focus their privacy concerns on invisible trackers and cookies while overlooking the semantic signals embedded in visible ad creatives (Yao et al., 2017). Similarly, app or plugin store reviews often focus on code safety rather than the privacy implications of data inferred from legitimate content access (Singh et al., 2025). This creates a regulatory blind spot where a benign-looking extension can silently harvest ad content without triggering security alarms. Furthermore, by leveraging the zero-shot capabilities of LLMs, the adversary can automate this process for scalable profiling. As we demonstrate in Section 5.3.1, actionable profiles can be reconstructed from short observation windows, meaning the adversary does not need to maintain long-term persistence to be effective. Ultimately, this vector offers a distinct strategic advantage: it enables off-platform profiling that evades the platform’s own privacy safeguards, such as the removal of sensitive targeting options. The adversary effectively exploits the optimisation logic of the ad platform to build sensitive user profiles at a low marginal cost and without leaving an audit trail.
@@ 4.1. Multimodal Ad Understanding
To transform raw, multimodal ad data into a structured representation suitable for profiling, we employ an LLM-driven extraction process. Let each advertisement $d_{i}$ in our dataset consist of visual content $V_{i}$ (one or more images) and textual metadata $T_{i}$ (title, body, call-to-action). We utilise Gemini 2.0 Flash (Google DeepMind, 2025) as our inference engine $\mathcal{M}$, selected for its balance of multimodal reasoning and cost-efficiency required to process $N>435,000$ impressions. For each ad, conditioned on an extraction prompt $\mathcal{P}_{\text{extract}}$, the model $\mathcal{M}$ maps the input into a structured feature set $F_{i}$: $F_{i}=\mathcal{M}(V_{i},T_{i}\mid\mathcal{P}_{\text{extract}})=\{c_{i},\mathcal{D}_{i},\mathcal{L}_{i},\mathcal{E}_{i}\}$.
This structured output $F_{i}$ encapsulates four semantic components: (1) a Caption ($c_{i}$) summarising the core message and visual elements of the ad; (2) Descriptive Categories ($\mathcal{D}_{i}$), a set of free-form labels characterising tone, style, and communication strategy; (3) IAB Categories ($\mathcal{L}_{i}$), a multi-label zero-shot classification into 45 IAB-defined categories (IAB Technology Laboratory, ); and (4) Key Entities ($\mathcal{E}_{i}$), a list of extracted specific brands, products, or locations. This transformation $d_{i}\to F_{i}$ extracts high-dimensional multimodal signals into a concise textual representation. We validate the quality of these model-generated features in Appendix C and discuss a case of Human–AI disagreement in Appendix D.
@@ 4.2. Session-Level User Profile Reconstruction
To capture short-term profiling signals, we segment the continuous user stream $\mathcal{S}_{u}$ into discrete sessions. Let a session $\mathcal{S}_{s}$ be a chronologically ordered sequence of $n$ structured ad features observed within a continuous browsing window (segmentation details in Section 5.2): $\mathcal{S}_{s}=(F_{s_{1}},F_{s_{2}},\dots,F_{s_{n}})$, where each $F_{s_{i}}$ corresponds to the structured feature set extracted in the previous stage. We construct a textual representation $T_{s}$ by concatenating the components of each $F\in\mathcal{S}_{s}$ in temporal order.